OpenStudy (anonymous):

Graph the circle with center at (3,-2), which also passes through the point (0,2)

OpenStudy (anonymous):

The first thing you have to do is use the distance formula to get the radius. That will be the distance from the center to the other point. Are you able to do that, or do you need help with that?

OpenStudy (anonymous):

need help, what is the formula

OpenStudy (anonymous):

btw, I was graphing it while you were answering, it looks like:

OpenStudy (anonymous):

\[d = \sqrt{(x _{2} - x _{1})^{2} + (y _{2} - y _{1})^{2}}\]

OpenStudy (anonymous):

That "d" will be your (r)adius for the equation of a circle with center at (h, k):\[(x - h)^{2} + (y - k)^{2} = r ^{2}\]

OpenStudy (anonymous):

so x2 is 0 and x1 is 3

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Are you able to get the radius?

OpenStudy (anonymous):

i got -3?

OpenStudy (anonymous):

Are you saying you got -3 for the radius?

OpenStudy (anonymous):

I got -3 as the answer to that formula

OpenStudy (anonymous):

or 1.414

OpenStudy (anonymous):

Well, if you did, you took the square root of something and got -3. It will help you immensely if you show me your work and I can then see where you went wrong.

OpenStudy (anonymous):

Actually, looking at those numbers, I can see where you probably went wrong without you even writing it out. 2 - (-2) = 4 You could also go: -2 - 2 = -4 Either one squared is 16

OpenStudy (anonymous):

\[r = d = \sqrt{(3 - 0)^{2} + (-2 - 2)^{2}} = \sqrt{9 + 16} = \sqrt{25} = 5\]

OpenStudy (anonymous):

(0-3)^2+(2+2)^2 ....-3^2=9 9+16=25 so 5?

OpenStudy (anonymous):

Yes, you can use either point for point #1 so either way works.

OpenStudy (anonymous):

So, the equation for the circle is:\[(x - 3)^{2} + (y + 2)^{2} = 25\]

OpenStudy (anonymous):

All good now, @adillie ?

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

uw! Good luck to you in all of your studies and thx for the recognition! @adillie

OpenStudy (anonymous):

I have one more question... is (3,3) (8,-2) (-2,-2) (3,-7) points on that circle?

OpenStudy (anonymous):

I just got back to my computer. Yes, they are all points on that circle. You can determine that from either the equation or the graph. @adillie